OpenStudy (anonymous):

A number is 8 more than 32. Find the number. Let n = number.

OpenStudy (anonymous):

n=(32)+8

OpenStudy (anonymous):

40

OpenStudy (anonymous):

32 + 8 = n 32 + 8 = 40 n = 40

OpenStudy (anonymous):

thankx